### v1

# A multi-band study of pulsar glitches with Fermi-LAT and Parkes

Abstract
Pulsar glitch is a phenomenon characterized by abrupt changes in the spin period over less than a minute. We present a comprehensive analysis of glitches in four gamma-ray pulsars by combining the timing observation data of Fermi Large Area Telescope (Fermi-LAT) and Parkes 64 m radio telescope. The timing data of five pulsars, namely PSRs J1028--5819, J1420--6048, J1509--5850, J1709--4429 (B1706--44) and J1718--3825, spanning over 14 years of observations for each, are examined. A total of 12 glitches are identified in four pulsars, including a previously unreported glitch. That is, a new small glitch is identified for PSR J1718--3825 in MJD $\sim$ 59121(8), and the fractional glitch size was $\Delta\nu/\nu\sim 1.9(2)\times 10^{-9}$. For PSR J1420--6048, our investigation confirms the existence of two linear recovery terms during the evolution of $\dot{\nu}$ subsequent to glitches 4, 6 and 8, and identified an exponential recovery process in glitch 8, with $Q=0.0131(5)$, $\tau_{\rm d}=100(6)$ d. Regarding the fourth glitch of PSR J1709--4429, our analysis reveals the presence of two exponential recovery terms with healing parameters and decay time-scales $Q$1=0.0104(5), $\tau_{\rm d1}=72(4)$ d and $Q$2 = 0.006(1), $\tau_{\rm d2}=4.2(6)$ d, respectively. For the remaining previously reported glitches, we refine the glitch epochs and glitch observables through precise fitting of the timing residual data. We extensively discuss how multi-band data of glitches can help better characterize the glitch recoveries and constrain the underlying physics of glitch events. We demonstrate that the accumulation of observational data reveals the rich complexity of the glitch phenomenon, aiding in the search for a well-established interpretation.
keywords:
stars: neutron – equation of state – pulsar: gamma-ray – pulsars: individual (PSR J1028--5819) – pulsars: individual (PSR J1420--6048) – pulsars: individual (PSR J1509--5850) – pulsars: individual (PSR J1709--4429 (B1706--44)) – pulsars: individual (PSR J1718--3825).
pubyear: 2024pagerange: A multi-band study of pulsar glitches with Fermi-LAT and ParkesLABEL:lastpage
1 Introduction
[…]
Table 1: Parameters and data spans for five gamma-ray pulsars studied in this paper.
PSR RA DEC $P$ $\dot{P}$ $B_{\rm s}$ Age DM Fermi-LAT span Parkes span
(hh:mm:ss) ($\ast^{\degr}:\ast^{\arcmin}:\ast^{\arcsec}$) (s) (10-15) ($10^{12}$ G) (kyr) (cm-3pc) (MJD) (MJD)
J1028--5819a 10:28:27.9(1) --58:19:06.21(6) 0.091403 16.1000 1.23 90.0 96.525 54687–59809 55233–59182
J1420--6048b 14:20:08.237(16) --60:48:16.43(15) 0.068180 83.1670 2.41 13.0 360.150 54722–59453 54303–59182
J1509--5850c 15:09:27.156(7) --58:50:56.01(8) 0.088925 9.1663 0.91 154.0 142.100 54713–59510 54303–59182
J1709--4429d 17:09:42.728(2) --44:29:08.24(6) 0.102459 92.9845 3.12 17.5 75.593 54684–59573 54303–59182
J1718--3825e 17:18:13.565(4) --38:25:18.06(15) 0.074670 13.2217 1.01 89.5 247.460 54741–59491 54303–59238
  • Note. References for parameters of these pulsars:
  • a Keith et al. (2008);
  • b D’Amico et al. (2001);
  • c Kramer et al. (2003);
  • d Johnston et al. (1992);
  • e Manchester et al. (2001).
Gamma-ray pulsars  [^1] Except millisecond pulsars, characterized as youthful neutron stars with high spin-down rates active in the gamma-ray band of emitted radiation, are good target sources for searching and studying glitches (Sokolova & Panin, 2022). More than 150 young gamma-ray pulsars have been reported so far, with more than 80 among them having exhibited radio emissions (Smith et al., 2023). Studying glitches in gamma-ray pulsars holds the potential to not only glean valuable insights into the internal structure of neutron stars but also to offer a window into the dynamic evolution of the magnetosphere (Kou et al., 2018).
Currently, there has been some progress in research focused on glitches in gamma-ray pulsars. PSR J2021+4026 is the only glitching pulsar for which a gamma-ray flux jump has been observed (Allafort et al., 2013; Takata et al., 2020). This phenomenon has been linked to glitches, with studies by Ng et al. (2016); Zhao et al. (2017) indicating a plausible association between gamma-ray flux jumps and substantial shifts in the pulse profile of PSR J2021+4026. In a similar manner, Lin et al. (2021) reported the detection of a new glitch in PSR J1420--6048 using Fermi-LAT 11 yr timing data and found that the gamma-ray pulse profile changed significantly within certain specific time ranges. Gügercinoğlu et al. (2022) identified 20 glitches in 4 gamma-ray pulsars, and made a relatively accurate prediction of the waiting time of the next glitch through the vortex creep model.
In this paper, we study the glitches of five gamma-ray pulsars by combining the timing observation data of the Fermi-LAT and Parkes telescopes in the past 14 yr. We have identified 12 glitches in total and detected a new small glitch at PSR J1718--3825. Besides, we update the epoch and other parameters of the previously reported 11 glitches.
The structure of this paper is as follows: In Section 2, we briefly introduce the timing observations of Fermi-LAT and Parkes 64 m radio telescope and give an overview of the glitch data processing procedure. One superfluid model, namely the vortex creep model, that is used in glitch fitting is introduced in Section 3, following which we relate the glitch observables to the dynamical and structural properties of the neutron star. The results and underlying physical process of pulsar glitches studied in the paper are detailed in Section 4. We then discuss our results more extensively in Section 5 and finally summarize our findings in Section 6.
[…]
PSR Int. Epoch $\nu$ $\dot{\nu}$ $\ddot{\nu}$ $N_{\rm ToA}$ $T$ Data span RMS
[…]
J1028--5819 0–1 56292 10.94024439333(3) --192.67239(3) 7.72(1) 380 8.8 54687–57898 1536
1– 58858 10.9399877351(3) --193.3893(6) 31.0(5) 221 5.2 57909 – 59809 8827
J1420--6048 3–4 54542 14.6625943767(3) --1782.52(1) 1130(18) 16 0.9 54303–54634 2
4–5 55279 14.661471259(2) --1781.60(3) 1549(12) 58 2.0 54652–55397 5256
5–6 55841 14.660623935(1) --1785.510(4) 1003(7) 53 2.3 55408–56256 3950
6–7 56704 14.659321169(2) --1785.782(9) 714(14) 51 2.6 56256–57212 9932
7–8 57890 14.657509511(3) --1783.284(6) 765(7) 62 3.6 57228–58555 17187
8– 59030 14.655774598(1) --1784.302(5) 983(7) 38 2.5 58557–59453 3805
J1709--4429 3–4 54496 9.7566104500(1) --883.9007(7) 330(3) 21 1.1 54303–54690 30
4–5 55516 9.7558562250(5) --885.766(1) 239.2(9) 435 4.5 54693–56337 17667
5–6 57256 9.7545531968(5) --885.721(1) 197.8(8) 493 5.0 56340–58173 21964
6– 58874 9.7533385441(9) --886.015(2) 229(2) 249 3.8 58173–59582 19600
J1718--3825 0–1 54605 13.3915736271(3) --236.412(1) 28(3) 24 2.2 54303–54920 513
1–2 56466 13.39119382383(4) --236.05539(7) 20.75(3) 106 8.2 54997–57940 104
2–3 58524 13.3907744744(1) --235.7702(3) 17.2(3) 45 3.2 57964–59113 533
3– 59301 13.3906162607(6) --235.65(1) -- 8 1.1 59128 – 59491 988
J1509--5850 56906 11.2452960325(2) --115.9029(5) -- 172 14.3 54303–59510 34250
2 Observation and analysis
In this work, we analysed the timing observation data of Fermi-LAT and Parkes 64 m radio telescope in Australia from 2007 to 2022 for PSRs J1028--5819, J1420--6048, J1509--5850, J1709--4429 and J1718--3825. Table 1 presents the detailed parameters of pulsars. The first three columns are the pulsar name (J2000) and the position [right ascension (RA) and declination (DEC)] in the celestial coordinate system. The fourth and fifth columns are the pulsar spin period ($P$) and its first time derivative ($\dot{P}$). The sixth and seventh columns are the inferred surface dipole magnetic field strength at the stellar equator ($B_{\rm s}=3.2\times 10^{19}\sqrt{P\dot{P}}$ G) and the characteristic age ($\tau_{\rm c}=P/2\dot{P}$) of pulsars. The last three columns are the pulsar’s dispersion measure (DM) and the timing observation data time span of the Fermi-LAT and the Parkes telescopes for the corresponding sources.
[…]
We also collected the timing observation data of these five gamma-ray pulsars from 2007 to 2020 on the Parkes pulsar data archive [^3] https://data.csiro.au/domain/atnf. Manchester et al. (2013) has introduced the timing observation of Parkes 64 m radio telescope in detail. Briefly, Parkes observation was carried out at 20 cm band, mainly using the multi-beam receiver, with a centre frequency of 1369 MHz and bandwidth of 256 MHz (Staveley-Smith et al., 1996; Hobbs et al., 2020). The Parkes Digital Filter Bank system used includes PDFB1, PDFB2, PDFB3 and PDFB4, and the generated file has 1024 frequency channels, each channel bandwidth is 0.25 MHz. The observation interval of each pulsar is 2–4 weeks, the observed sub-integration time was 30 s, and the integration time was 2–15 mins.
For radio timing data, PSRCHIVE was used to eliminate the dispersion effects and radio frequency interference (RFI), while superimposing to obtain an average integrated pulse profile for each observation (Hotan et al., 2004; van Straten et al., 2012). The timing data were partitioned based on glitch epoch, and then PSRADD was used to superimpose the average pulse profile of each observation in the polarization, time domain and frequency to obtain the corresponding total integrated pulse profile. The total integrated pulse profile with the highest signal-to-noise ratio (S/N) was selected as the standard pulse profile. The PAT tool within PSRCHIVE was then used to cross-correlate each mean profile with the standard profile, obtaining highly accurate pulse times of arrival (ToAs; Taylor 1992). Finally, the Jet Propulsion Laboratory planetary ephemeris DE440 (Park et al., 2021) and the Barycentric Coordinate Time (TCB) were used to convert these local ToAs to SSB, thereby eliminating the effect of the Earth motion on the ToAs.
We combined the gamma-ray and radio ToAs, and used the high-precision timing analysis software packages TEMPO2 (Hobbs et al., 2006; Edwards et al., 2006) to achieve the timing solution. We introduced two parameters, EFAC and EQUAD, to correct the uncertainty of each ToA acquired from different telescopes and terminals (Wang et al., 2015). Among them, EFAC is a free parameter that adjusts unaccounted instrumental errors, and EQUAD accounts for additional time-independent white noise processes (Lentati et al., 2014). The values of these two parameters were determined by the EFACEQUAD plug-in of TEMPO2, and this method did not involve Bayesian techniques. After correction, the relationship between the initial uncertainty ($\delta$) of ToAs and its new value (${\delta}_{\rm s}$) was defined as ${\delta}_{\rm s}^{\rm 2}={\rm EFAC}^{2}\times(\delta^{2}+{\rm EQUAD}^{2})$. After adding the error correction of ToA, the reduced chi-square values of each fitting region are distributed between 0.95–1.05.
The pulsar position is one important factor that affects the precision of glitch parameter, but for some sources, the position obtained from PSRCAT is not accurate enough. For these pulsars, we used the "Cholesky" method to redefine the position of the pulsars (Coles et al., 2011). To do this, we selected a sufficiently long length of data that not include any glitch, invoked the TEMPO2 plug-in SPECTRALMODEL to fit the red noise model (power law model of red noise: $P(f)={A}{[1+(f/f_{\rm c})^{2}]^{(-\alpha/2)}}$, where $A$, $\alpha$, and $f_{\rm c}$ are amplitude, spectral index, and corner frequency, respectively), then referred to this model to obtain the accurate position. After completing the above steps, we employed TEMPO2 to fit ToAs to obtain the best rotation phase model of pulsars. By subtracting the actual observed ToAs from the ToAs predicted by the model, timing residuals were obtained.
According to the pulsar timing model, the evolution of pulse phase with time, $\phi(t)$, can be expressed as (Edwards et al., 2006):
$\phi(t)=\phi_{\rm 0}+\nu(t-t_{\rm 0})+\frac{\dot{\nu}}{2!}(t-t_{\rm 0})^{2}+\frac{\ddot{\nu}}{3!}(t-t_{\rm 0})^{3}\ ,$ (1)
where $t_{\rm 0}$ represent certain initial epoch, and $\phi_{\rm 0}$ is the initial pulse phase at $t=t_{\rm 0}$. Here, $\nu$ is the spin frequency of pulsar, $\dot{\nu}$ and $\ddot{\nu}$ represent the first and second derivatives of the frequency, respectively.
Timing residuals were obtained by comparing observations with model-predicted ToAs. We can use TEMPO2 to visualize timing residuals. In an ideal state, the timing residuals should be randomly distributed around zero. If, upon visual inspection, the residuals appear discontinuous or show significant deflection, it may indicate the occurrence of a glitch event in the pulsar. Assuming a glitch event occurs at epoch $t_{\rm g}$, it may result in a change in the pulse phase is $\Delta\phi$, a permanent increase in the spin frequency, $\Delta\nu_{\rm p}$, and a permanent increase in its first derivative $\Delta\dot{\nu}_{\rm p}$. It should be noted that the post-glitch recovery process typically involves one or more exponential decay processes. If the frequency change during the exponential recovery process of the $i$-th phase is $\Delta\nu_{\rm d_{i}}$, and the corresponding timescale is $\tau_{\rm d_{i}}$, we can describe the change of post-glitch pulse phase according to the above-mentioned relevant glitch parameters as (Edwards et al., 2006; Yu et al., 2013):
$\begin{split}\phi_{\rm g}=&\Delta\phi+\Delta\nu_{\rm p}(t-t_{\rm g})+\frac{1}{2}\Delta\dot{\nu}_{\rm p}(t-t_{\rm g})^{2}\\ &+\sum_{\rm i}{\Delta\nu_{\rm d_{i}}\tau_{\rm d_{i}}[1-e^{-(t-t_{\rm g})/\tau_{\rm d_{i}}}]}\ .\end{split}$ (2)
According to the above equation, the total fractional glitch size and its first derivative are respectively represented as:
$\frac{\Delta{\nu}}{{\nu}}=\frac{\Delta{\nu}_{\rm p}+\sum_{\rm i}\Delta\nu_{\rm d_{i}}}{\nu}\ ,$ (3)
and
$\frac{\Delta\dot{\nu}}{\dot{\nu}}=\frac{\Delta\dot{\nu}_{\rm p}-\sum_{\rm i}{\Delta\nu_{\rm d_{i}}/\tau_{\rm d_{i}}}}{\dot{\nu}}\ .$ (4)
Finally, the parameter $Q$ that quantifies the fraction of glitch recovery can be defined as:
$Q=\frac{\sum_{\rm i}{\Delta\nu_{\rm d_{i}}}}{\sum_{\rm i}{\Delta\nu_{\rm d_{i}}}+\Delta\nu_{\rm p}}\ .$ (5)
Table 3: Glitch parameters obtained through fitting methods based on Eqs. (1) and (2). See text for details.
PSR Gl. No. Epoch $\Delta\nu/\nu$ $\Delta\dot{\nu}/\dot{\nu}$ NEW? $Q$ $\tau_{\rm d}$ RMS Data span
[…]
J1028--5819 1 57904(6) 2284.0(9) 20(2) P 0.0060(2) 62(7) 636 57358–58243
J1420--6048 4 54653(19) 937.6(9) 6.4(3) P 0.0119(6) 53(7) 3 54310–54949
5 55400(9) 1367(15) 5.3(8) P 0.016(8) 187(82) 991 55040–55804
6 56256(1) 1967(2) 8.6(9) P 0.0099(7) 44(8) 818 55956–56536
7 57216(12) 1206(1) 4.9(2) P 0.0183(6) 102(7) 943 56800–58110
8 58555(2) 1480(1) 6.0(1) P 0.0131(5) 100(6) 434 58141–59002
J1709--4429 4 54691(2) 2777(4) 64(15) P 0.0104(5) 72(4) 387 55415–55167
0.006(1) 4.2(6)
5 56339(2) 2964(4) 9.7(8) P 0.0095(8) 61(6) 747 56158–56603
6 58175(2) 2438(3) 10(1) P 0.0087(7) 47(8) 1189 57918–58401
J1718--3825 1 54952(44) 1.98(6) --0.10(3) P 50 54289–55860
2 57950(14) 7.37(9) 0.21(6) P 493 57160–58515
3 59121(8) 1.9(2) --0.12(9) Y 596 58510–59527
Table 4: Vortex creep model fit results for 9 large glitches in three gamma-ray pulsars analysed in this work: The characteristic relaxation time of exponential decay $\tau_{\rm creep,exp}$ from the linear response regime in vortex creep model can be confronted with that of the TEMPO2 fitting $\tau_{\rm d}$ listed in Table 3. $\tau_{\rm nl}$ ($\tau_{\rm s}$ in the case of single non-linear regime) is the non-linear creep relaxation time. Post-glitch recovery is expected to end when the relaxations of the non-linear creep regions are over, and glitch induced changes in the spin-down rate recover completely after a waiting time $t_{0}$. Thus the theoretical expectation $t_{\rm th}=\mbox{max}(t_{0},t_{\rm single})$, i.e., the maximum of Eq. (11) and Eq. (13) gives a prediction for the inter-glitch time.
PSR Gl. No. $\tau_{\rm creep,exp}$ $\tau_{\rm nl}$ $t_{0}$ $\tau_{\rm s}$ $t_{\rm s}$ $t_{\rm th}$ $t_{\rm obs}$
(d) (d) (d) (d) (d) (d) (d)
J1028--5819 1 96(12) 71 4890(269) - - 4890(269) -
J1420--6048 4 41(11) - - 116(10) 349(8) 698(32) 747(28)
5 - 109(23) 125(42) 109(6) 638(9) 964(19) 856(10)
6 84(11) 20(12) 562(13) 98(11) 829(6) 1124(32) 960(13)
7 109(32) 779(554) 938(425) - - 938(425) 1339(14)
8 86(39) 19 823(16) 26(11) 623(11) 823(16) -
J1709--4429 4 86(9) 3.4 1680(16) - - 1680(16) 1648(4)
5 95(19) 28 1861(25) - - 1861(25) 1836(4)
6 93(9) 3.2 1850(26) - - 1850(26) -
3 Glitch trigger and post-glitch relaxation in the vortex creep model
The observations of pulsar glitches provided the first evidence of superfluidity in neutron-star interiors (Baym et al., 1969). Previously, our works focused on the amplitude of the glitch and explored whether modern equation of state (EOS) of pulsars can reconcile with the available data (Li, 2015; Li et al., 2016), mainly in the framework of the two-component model (Baym et al., 1969; Anderson & Itoh, 1975), i.e., the superfluid component and the other is the charged component. With various post-glitch recovery data available (Ge et al., 2020; Gügercinoğlu et al., 2022; Liu et al., 2024; Tuo et al., 2024), we here adopt the vortex creep model and confront it with both the observed amplitudes and post-glitch decays (Shang & Li, 2021), in our endeavor to search for a firmly grounded paradigm of glitch phenomenon.
The vortex creep model (Alpar et al., 1984, 1989) aimed at explaining both the glitch triggering and the post-glitch relaxation. In the theory, there exist a number of distinct superfluid regions with different pinning energies in the inner crust. The neutron superfluid vortices are pinned to nuclei or interstitial positions, developing a lag $\omega=\Omega_{\mathrm{s}}-\Omega_{\mathrm{c}}$ between the pinned neutron superfluid and the crustal component, which spins down under the external pulsar braking torque. Here $\Omega_{\mathrm{s}}$ and $\Omega_{\rm c}$ are the superfluid and the crustal angular velocities, respectively. The radially outward motion of vortices caused by thermal excitation overcoming pinning barriers is termed vortex creep. The continuous vortex creep results in angular momentum transfer from the crustal superfluid to the crust and the coupling between them, a steady state $\omega_{\infty}$ can reached when the crustal superfluid and crust share the same spin-down rate $\dot{\Omega}_{\mathrm{s}}=\dot{\Omega}_{\rm c}=\dot{\Omega}_{\infty}$. In the steady state, the star’s spin-down appears as if an external torque $N_{\rm{ext}}$ acts on the star’s total moment of inertia $I$, $N_{\rm{ext}}=I\dot{\Omega}_{\infty}$.
A glitch occurs if a local fluctuation raise $\omega$ to above $\omega_{\mathrm{cr}}$, here $\omega_{\mathrm{cr}}$ is a critical lag where the Magnus force exceeds the pinning force and it depends on the region where the vortices are located. When a glitch occurs, a number of vortices unpin suddenly from pinning sites and move outward; at the same time, the steady state is broken down with a perturbation $\delta\omega$ deviated from $\omega_{\infty}$. Due to the conservation of angular momentum, the angular velocity jump $\Delta\Omega_{\mathrm{c}}$ for the observed crust is
$I_{\mathrm{c}}\Delta\Omega_{c}=\sum_{i}I_{\rm i}\delta\Omega_{\rm i},$ (6)
where $I_{\rm i}$ are the moment of inertia of distinct superfluid regions and $\delta\Omega_{\rm i}$ are the corresponding decrease in the angular velocity. The derivation $\delta\omega$ in a superfluid region can be expressed by $\delta\omega_{\rm i}=\Delta\Omega_{\mathrm{c}}+\delta\Omega_{\rm i}$. After a glitch, the star returns to a new steady state, towards original pre-glitch state, as a result of the recovery of vortex creep. Depending on the temperature and the pinning properties, the vortex creep can have a linear or non-linear response to the perturbation $\delta\omega$ in the different superfluid regions (Alpar et al., 1989).
In the linear response regime, approach to the steady state has a linear dependence on the initial perturbation $\delta\omega(0)$. The post-glitch recovery in this regime can be written by
$\Delta\dot{\Omega}_{\mathrm{c,i}}(t)=-\frac{I_{\rm i}}{I}\frac{\delta\omega_{\rm i}(0)\rm{e}^{-t/\tau_{\mathrm{l,i}}}}{\tau_{\mathrm{l,i}}},$ (7)
where $\tau_{\mathrm{l,i}}$ is the relaxation time of this regime and $\delta\omega_{\rm i}(0)$ is the value of $\delta\omega_{\rm i}$ at the time when a glitch occurs. The relaxation behaviour is exponential decay with the characteristic relaxation time
$\tau_{\mathrm{l}}=\frac{kT}{E_{\mathrm{p}}}\frac{\omega_{\mathrm{cr}}r}{4\Omega_{\rm c}v_{0}}\exp(\frac{E_{\mathrm{p}}}{kT}),$ (8)
where $kT$ and $E_{\mathrm{p}}$ are the internal temperature and the pinning energy in this regime; $r$ is the distance of this regime from the spin axis; $v_{0}\approx 10^{5}$$10^{7}$ cm s-1 is the typical velocity of microscopic vortex motion (Gügercinoğlu & Alpar, 2016; Gügercinoğlu et al., 2023). In the outer core, vortex creep against flux tubes gives a similar post-glitch exponential relaxation with the relaxation time (Gügercinoğlu & Alpar, 2014),
[…]
where $\dot{\Omega}_{\mathrm{c}}$ is the spin-down rate, $R$ is the radius of the outer core, $x_{\rm p}$ is the proton fraction in the neutron star core, $\rho$ is the matter density, $m_{\rm p}^{*}(m_{\rm p})$ is the effective (bare) mass of protons, and $B_{\phi}$ is the toroidal component of the magnetic field strength. A linear regime occurs when the temperature is high enough compared to the pinning energy, hence we have $\omega_{\infty}\ll\omega_{\mathrm{cr}}$ and $\delta\omega_{\rm i}(0)\approx\Delta\Omega_{\mathrm{c}}$.
In the non-linear response regime, at the other extreme, approach to the steady state depends non-linearly on the initial perturbation $\delta\omega(0)$. For a single crustal superfluid region, $k$, in the non-linear regime, the post-glitch response is
$\frac{\Delta\dot{\Omega}_{\mathrm{c},k}(t)}{\dot{\Omega}_{\mathrm{c}}}=-\frac{I_{\rm s,k}}{I}\left\{1-\frac{1}{1+[\rm{e}^{(t_{\mathrm{0},k}/\tau_{\mathrm{n},k})}-1]\rm{e}^{(-t/\tau_{\mathrm{n},k})}}\right\}\ ,$ (10)
where
$t_{\mathrm{0,k}}=\delta\omega_{\rm k}(0)/|\dot{\Omega}|_{\infty}\ ,$ (11)
and
$\tau_{\mathrm{n}}=\frac{kT}{E_{\mathrm{p}}}\frac{\omega_{\mathrm{cr}}}{|\dot{\Omega}|_{\infty}}\ .$ (12)
Any observed value of $|\dot{\Omega}_{\mathrm{c}}|$ can represent $|\dot{\Omega}|_{\infty}$ after considering the small magnitude $\Delta\dot{\Omega}_{\mathrm{c}}/\dot{\Omega}_{\mathrm{c}}\sim 10^{-2}$. In the non-linear regime, the interior temperature is sufficiently low compared to the pinning energy, hence we have $\omega_{\infty}\lesssim\omega_{\mathrm{cr}}$ and $\delta\omega_{\rm k}(0)\approx\delta\Omega_{\rm s,k}$. Eq. (10) has the form of making a “shoulder” such that at about $t_{\mathrm{0,k}}$ after the glitch, the spin-down rate decreases rapidly by $I_{\rm s,k}/I$ in a time $\sim 3\tau_{\mathrm{n,k}}$ quasi exponentially. This response resembles the behaviour of the Fermi distribution function of statistical mechanics for non-zero temperatures (Alpar et al., 1984). Single superfluid region’s recovery is thus completed on a timescale,
$t_{\rm{single,k},}=t_{\rm 0,k}+3\tau_{\mathrm{n,k}}.$ (13)
In several neighbouring single non-linear regime regions, we can integrate explicitly the contributions from these single non-linear regime regions by making the assumption that the post-glitch superfluid angular velocity decreases linearly in $r$ from $\delta\Omega_{\mathrm{A}}$ over the region (Alpar et al., 1984), then the post-glitch response is (Alpar et al., 1996):
$\frac{\Delta\dot{\Omega}_{\mathrm{c,A}}(t)}{\dot{\Omega}_{\mathrm{c}}}=\frac{I_{\mathrm{A}}}{I}\left\{1-\frac{1-(\tau_{\mathrm{n,A}}/t_{\mathrm{0,A}})\ln\left[1+(\mathrm{e}^{t_{\mathrm{0,A}}/\tau_{\mathrm{n,A}}}-1)\mathrm{e}^{-\frac{t}{\tau_{\mathrm{n,A}}}}\right]}{1-\mathrm{e}^{-\frac{t}{\tau_{\mathrm{n,A}}}}}\right\},$ (14)
Here, the total moment of inertia of the non-linear creep regions affected by vortex unpinning events is denoted by $I_{\mathrm{A}}$. Gügercinoğlu & Alpar (2020) obtained the following estimate for $\tau_{\rm{n}}$ by incorporating density dependence of the pinning energy throughout the crustal superfluid and temperature evolution of the neutron star,
[…]
where $\tau_{\rm sd}=\nu/(2\dot{\nu})$ is the spin-down timescale (characteristic age) of the pulsar. In the limit of $t_{\mathrm{0,A}}\gg\tau_{\mathrm{n,A}}$, Eq. (14) reduces to
$\frac{\Delta\dot{\Omega}_{\mathrm{c,A}}(t)}{\dot{\Omega}_{\mathrm{c}}}=\frac{I_{\mathrm{A}}}{I}\left(1-\frac{t}{t_{\mathrm{0,A}}}\right),$ (16)
with a constant, anomalously large $\ddot{\Omega}_{\mathrm{c}}=(I_{\rm A}/I)(|\dot{\Omega}_{\rm c}|/t_{\rm 0,A})$ (Alpar & Baykal, 2006). The post-glitch recovery is expected to come to end when the relaxations of the non-linear creep regions are over. So the theoretical expectation $t_{\rm th}=\mbox{max}(t_{\mathrm{0,A}},t_{\rm single})$, i.e. maximum of Eq. (11) and Eq. (13) gives a prediction for the inter-glitch time. In practice, we usually choose a summation of several post-glitch recoveries of linear and non-linear regions to fit the post-glitch observations, e.g. a linear region with $I_{\rm{exp}}$ and $\tau_{\rm{creep,exp}}$, a single non-linear region with $I_{\rm{s}}$, $t_{\rm{s}}$ and $\tau_{\rm{s}}$, and a continuous nonlinear regions with $I_{\rm{A}}$, $t_{0}$, and $\tau_{\rm{nl}}$. $\Delta\dot{\Omega}_{\mathrm{c}}/\dot{\Omega}_{\mathrm{c}}$ is related to $\Delta\dot{\nu}/\dot{\nu}$ by $\Omega=2\pi\nu$.
Table 5: Moment of inertia of distinct superfluid regions for 9 large glitches in three gamma-ray pulsars analysed in this work: $I_{\rm A}$ is the total moment of inertia of the non-linear creep regions affected by vortex unpinning events, i.e., $I_{\rm A}/I$ is the fractional moment of inertia of the superfluid regions that gave rise to the glitch by collective unpinning event; the vortex then move to other trapped region by crossing vortex-free region B with moment of inertia $I_{\rm B}$, i.e., $I_{\rm B}/I$ is the fractional moment of inertia of the superfluid regions within which unpinned vortices moved; $I_{\rm exp}/I$ is fractional moment of inertia of the exponentially decaying superfluid region. $I_{\rm cs}/I$ is the fractional moment of inertia of the crustal superfluid that involved in the corresponding glitch event, which corrected for crustal entrainment effect average enhancement factor $\langle m_{\rm n}^{*}/m_{\rm n}\rangle=5$ (Delsate et al., 2016).
PSR Gl. No. $I_{\rm exp}/I$ $I_{\rm A}/I$ $I_{\rm s}/I$ $I_{\rm B}/I$ $I_{\rm cs}/I$
[…]
J1028--5819 1 7.92(1.63) 5.25(8) - 2.83(16) 3.35(16)
J1420--6048 4 15.34(6.41) - 4.26(31) 2.13(6) 2.56(6)
5 - 7.03(3.21) 2.72(13) 1.36(10) 2.34(34)
6 15.05(2.52) 1.47(17) 1.92(7) 1.20(1) 1.54(2)
7 16.45(5.56) 16.44(1.73) - 0.96(40) 2.05(1.98)
8 8.70(5.95) 3.13(16) 0.45(6) 0.87(2) 1.23(2)
J1709--4429 4 14.35(26) 3.66(4) - 1.93(2) 2.30(2)
5 14.03(7.12) 3.37(6) - 1.86(3) 2.20(3)
6 16.67(2.44) 3.36(5) - 1.51(2) 1.85(2)
During the glitch, the vortices unpinned from the pinning sites of nonlinear creep region A, then move outward to the other nonlinear creep region A by crossing through region B with the moment of inertia $I_{\mathrm{B}}$. Region B contributes to the angular momentum transfer at glitches while it cannot sustain the pinned vortices and hence cannot participate the spin-down caused by vortex creep (Cheng et al., 1988). According to the Eq. (6), $I_{\mathrm{B}}$ can be determined by
$I_{\mathrm{c}}\Delta\Omega_{\mathrm{c}}=(I_{\mathrm{A}}/2+I_{\mathrm{B}}+I_{\mathrm{s}})\delta\Omega_{\mathrm{s}},$ (17)
where the factor 1/2 accounts for linear decrease of $\delta\Omega_{\mathrm{A}}$ within $I_{\mathrm{A}}$. The total crustal superfluid moment of inertia involved a glitch is $I_{\rm{cs}}=I_{\rm{A}}+I_{\rm{s}}+I_{\rm{B}}$. Importantly, it has been argued that “entrainment” of the neutron superfluid by the crystalline structure of the crust greatly reduces its mobility, reducing the superfluid angular momentum reservoir to be tapped at glitches (Andersson et al., 2012; Chamel, 2013). Due to this effect, moments of inertia of the superfluid components $I_{\rm A}$ and $I_{\rm s}$ obtained from the post-glitch timing fits should be multiplied by the same enhancement factor.  [^4] However, the mobility of superfluid neturons related to the effective neutron mass $m_{\rm n}^{*}/m_{\rm n}$ is an open problem which has been discussed actively in the literature, see e.g., Chamel (2005, 2012); Watanabe & Pethick (2017); Chamel (2017). More recently, it has been shown that in the so-called pasta phase, where sizeable crustal moment of inertia resides, the effect is reversed so that anti-entrainment with smaller than unity for $m_{\rm n}^{*}/m_{\rm n}$ realizes (Yoshimura & Sekizawa, 2024). Regularity of glitches and data accumulated over the years have important implications on the parts of the neutron star that drive the spin-up events and thereby the moments of inertia of the corresponding regions. Thus, the long-term glitch observations can be utilized to bring constraints on the EOS.
4 RESULTS
[…]
Table 3 reports the parameters of all the detected glitches. The first three columns are the pulsar name, the reference number of each glitch, and the glitch epoch. Glitch epoch is obtained by extrapolation where the phase continuity was lost in the timing solution. It is necessary to use Eq. (1) to fit the timing solutions of pre-glitch and post-glitch respectively. Glitch epoch is the middle of interval between the last pre-glitch observation time and the first post-glitch observation time. The uncertainty is half of the interval between the two observations. The other parameters are obtained through phase-coherent timing fits, that is, fitting Eq. (2) to ToAs, while the associated uncertainty is calculated using the error transfer equation. The fourth and fifth columns show the relative change in the spin frequency and the first derivative of frequency, respectively. In the sixth column, Y represents the new detected glitch in this work and P is the previously published glitch event. If significant exponential recovery process is observed, we will add exponential decay term to the TEMPO2 fitting, the results of these parameters are shown in the seventh and eighth columns. The last two columns are the RMS and the data span. The results of the timing analysis are shown in Figs. (1), (2), (3), (5) and (6), where value of each $\nu$ and $\dot{\nu}$ was obtained through independent fitting of ToAs of 150–600 d. In this work, we use the least squares method to fit the exponential decay and linear recovery of the spin-down rate after the glitch, respectively. Specifically, the red curves in these figures represent the fitting results of exponential recovery, while the blue and magenta dotted curves represent the fitting results of two different linear recoveries.
Next, we will specifically describe the glitch events. The fits within the framework of the vortex creep model (see Sec. 3) are done for 9 large glitches in our sample, using Eqs. (7), (10) and (14). The results are reported along with Figs. (1), (2), (3), (5) and in Table 4. One sees that the recovery fit from a different and physically-motivated functional form describes the observations equally well. Various moments of inertia from the vortex-creep-model fit are collected in Table 5. The results suggest that glitches involve only the crustal superfluid and possibly a portion of the outer core, with the adopted creep model providing sufficient capacity for the observed glitch amplitudes. In the next section of Sec. 5, we also show our results for the oscillations in the spin-down rates of PSRs J1509--5850 and J1718--3825, and reassess them within the vortex bending model. Moreover, beyond the common magnetic dipole-radiation scenario, we discuss the effect of the internal torque from bending vortex and inter-glitch creep contributions on the braking index.
4.1 PSR J1420--6048
Figure 1: Glitches in PSR J1420--6048: The vertical gray dashed line and the number of parentheses at the top represent the glitch epoch and the reference number of glitch occurrences, respectively. Panel (a) shows the relative change ($\Delta\nu$) of spin frequency, that is, the change in pulsar rotation frequency relative to the frequency model before the glitch. Panel (b) is an enlarged view of panel (a), showing the evolution details of $\Delta\nu$ after the glitch, which is obtained by subtracting its average value from $\Delta\nu$ of post-glitch. Panel (c) shows the evolution of the spin-down rate ($\dot{\nu}$) over time before and after the glitch. The red, blue, and magenta dashed lines represent the one exponential and two different linear recovery fitting results, respectively. In Panel (d), the green curves represent the fit using the vortex creep model (see Sec. 3), with the corresponding fit parameters provided in Table 4.
PSR J1420--6048 is an energetic Vela-like pulsar, first discovered during the Parkes multilateral pulsar survey by D’Amico et al. (2001). In addition to being observed in the radio band, this young pulsar has been detected in both the X-ray and gamma-ray bands (Weltevrede et al., 2010). X-ray pulsations (AX J1420.1--6049) were observed by Roberts et al. (2001) using ASCA, while gamma-ray emissions (GeV J1417--6100/3EG J1420--6038) were detected by the EGRET telescope onboard the Compton Gamma Ray Observatory (Nishida et al., 2003).
PSR J1420--6048 has a spin period of 68 ms and a period derivative of $\dot{P}\sim 83.167\times 10^{-15}$ (D’Amico et al., 2001). From these values, the pulsar’s surface magnetic field is estimated to be $B\sim 2.41\times 10^{12}$ G, and its characteristic age is $\sim$ 13 kyr (D’Amico et al., 2001). This pulsar also has a high spin-down energy loss rate of $\dot{E}\sim 1.0\times 10^{37}$ erg s-1, placing it among the 18 most energetic known pulsars. Given its youth and large energy loss rate, PSR J1420--6048 is a valued source for studying glitch activity.
To date, eight glitches have been observed in PSR J1420--6048 (Yu et al., 2013; Liu et al., 2021; Lower et al., 2021; Lin et al., 2021), with fractional glitch sizes on the order of $\Delta\nu/\nu\sim 10^{-6}$. These glitches generally exhibit exponential recoveries, with glitch intervals of 2–3 yr. Lin et al. (2021) found that in the gamma-ray band, the pulse profile of PSR J1420--6048 may change significantly over different time spans.
We have combined Fermi-LAT data with Parkes radio timing data to update the glitch epochs and parameters for glitches 4, 5, 6, 7, and 8, as shown in Table 3. Fig. 1 shows the 14 yr evolution of $\nu$ and $\dot{\nu}$ for PSR J1420--6048. Previous studies indicated no exponential decay component in the post-glitch recovery of glitch 8 (Lin et al., 2021). However, our analysis of the fitted timing residuals reveals a clear exponential recovery component, as shown in Fig. 1. We modeled the post-glitch recovery for glitch 8, finding an exponential timescale of 100(6) d and $Q=0.0131(5)$. Additionally, we updated the fractional glitch size and its first derivative to $\Delta\nu/\nu\sim 1480(1)\times 10^{-9}$ and $\Delta\dot{\nu}/\dot{\nu}\sim 6.0(1)\times 10^{-3}$, respectively.
From panel (c) of Fig. 1, we observe that glitches 4, 6, and 8 each exhibit two distinct linear recovery phases following an initial exponential recovery. This is particularly evident for glitch 4. To verify this, we performed exponential and linear fits to the evolution of $\dot{\nu}$. In panel (c), the red curve represents the exponential fit, while the blue and magenta dashed lines correspond to the first and second linear recovery phases, respectively. Just as shown in panel (c), the linear processes of glitch 4 and glitches 6, 8 are clearly different. For glitch 4, the slope of the first linear recovery phase ($\ddot{\nu}\sim 1.778\times 10^{-21}$ s-3) is nearly double that of the second phase ($\ddot{\nu}\sim 8.469\times 10^{-22}$ s-3). In contrast, glitches 6 and 8 show the opposite trend: the slopes of the second linear phase are larger than the first. Specifically, glitch 6 has $\ddot{\nu}$ values of $5.454\times 10^{-22}$ s-3 and $8.364\times 10^{-22}$ s-3, while glitch 8 exhibits values of $8.388\times 10^{-22}$ s-3 and $1.237\times 10^{-21}$ s-3 for the two linear phases, respectively. For glitches 5 and 7, a single linear recovery phase adequately describes the post-glitch evolution, with slopes of $1.185\times 10^{-21}$ s-3 and $6.327\times 10^{-22}$ s-3, respectively. This behaviour of two consecutive linear recovery processes in the post-glitch has not been previously reported for PSR J1420--6048, likely due to the limited timing data available from individual Parkes and Fermi-LAT observations.  [^6] Nevertheless, it should be mentioned here that the “shoulder” or “elongated-S” shape seen after the glitches in PSR J1420--6048 may also be due to an unaccounted part of the post-glitch recovery times which could be originated from unmodelled timing noise as discussed in (Antonopoulou et al., 2022). More discussions on recovery behaviour can be found below in Sec. 5.
We apply the vortex creep model to fit the five glitches of PSR J1420--6048, and the fit results are shown in Fig. 1, panel (d):
  • G4: We fit the fourth glitch (G4) of PSR J1420--6048 with Eqs. (7) and (10). As seen in Table 4, the vortex creep model prediction for the time to the next glitch from G4 is $t_{\rm th}=698(32)$ d which is in qualitative agreement with the observed inter-glitch time $t_{\rm obs}=747(28)$ d within uncertainties. The exponential decay time-scale obtained from the vortex creep model application $\tau_{\rm creep,exp}=41(11)$ days is slightly different than what found from fit with Eq. (2): $\tau_{\rm d}=53(7)$ d (see Table 3). The reason is that Eq. (10) has also an exponential dependence and dominates over the initial days of the data. The recoupling time-scale for the crustal superfluid associated with G4 is $\tau_{\rm s}=116(10)$ d. The total fractional crustal superfluid moment of inertia involved in G4 is $I_{\rm cs}/I=2.56(6)\%$.
  • G5: We fit the fifth glitch (G5) of PSR J1420--6048 with Eqs. (10) and (14). Due to the initial exponential dependencies of the two equations, there is no need to include a separate exponential function to fit the data of G5. As seen in Table 4, the vortex creep model prediction for the time to the next glitch from G5 is $t_{\rm th}=964(19)$ d, while the observed inter-glitch time $t_{\rm obs}=856(10)$ d. The origin of the discrepancy may lie in the fact that crustquake, which played the role of starting the vortex unpinning avalanche, probably left behind a tiny persistent shift in the spin-down rate by leading to a horizontal displacement of the broken platelet. Such effects on the vortex creep model estimates are also visible for some of the Crab (Gügercinoǧlu & Alpar, 2019) and the Vela (Akbal et al., 2017) glitches. The recoupling time-scale for the crustal superfluid associated with G5 is $\tau_{\rm s}=109(6)$ d. The total fractional crustal superfluid moment of inertia involved in G5 is $I_{\rm cs}/I=2.34(34)\%$. Both $\tau_{\rm s}$ and $I_{\rm cs}/I$ are quite similar for G4 and G5, suggesting that these two glitches have originated from the same crustal superfluid region.
  • G6: We fit the sixth glitch (G6) of PSR J1420--6048 with Eqs. (7), (10) and (14). As seen in Table 4, the exponential decay time-scale obtained from the vortex creep model application $\tau_{\rm creep,exp}=84(11)$ d is slightly different than $\tau_{\rm d}=44(8)$ d found from fit with Eq. (2) (see Table 3) again due to the initial exponential dependencies of Eqs. (10) and (14). The vortex creep model prediction for the time to the next glitch from G6 is $t_{\rm th}=1124(32)$ d, while the observed inter-glitch time is $t_{\rm obs}=960(13)$ d. However, as can be clearly seen from Fig. 1, the seventh glitch (G7) arrived before the recovery of G6 has been fully completed. The recoupling time-scale for the crustal superfluid associated with G6 is $\tau_{\rm s}=98(11)$ d, similar to the value of the previous two events. The total fractional crustal superfluid moment of inertia involved in G6 is $I_{\rm cs}/I=1.54(2)\%$.
  • G7: We fit the seventh glitch (G7) of PSR J1420--6048 with Eqs. (7), and (14). As seen in Table 4, the exponential decay time-scale obtained from the vortex creep model application $\tau_{\rm creep,exp}=109(32)$ d is consistent with $\tau_{\rm d}=102(7)$ d found from fit with Eq. (2) (see Table 3). As can be clearly seen from Fig. 1, the post-glitch spin-down rate of G7 over-recoveries, i.e. extends over the original pre-glitch level. One possible explanation of this peculiar and rare type of relaxation in terms of the vortex creep model is as follows. According to the vortex creep model, there are two nonlinear creep regions responsible for the post-glitch rotation and spin-down rate evolution, namely vortex traps (initiating a glitch through collective vortex unpinning) with moment of inertia $I_{\rm A}$ and vortex capacitor regions (inside of which vortices cannot creep but other freed lines may pass through of it and take part in angular momentum conservation) with moment of inertia $I_{\rm B}$. If some of these capacitor regions become activated such that vortices inside that region start to creep after their interaction with the unpinned ones, then with the excess angular momentum transfer of such activated vortices the post-glitch spin-down rate will extend over the pre-glitch level. The vortex creep model prediction for the time to the next glitch from G7 is $t_{\rm th}=938(425)$ d and agrees with the observed inter-glitch time $t_{\rm obs}=1339(14)$ d within errors. The recoupling time-scale for the crustal superfluid associated with G6 is $\tau_{\rm nl}=779(554)$ d and due to its being longer has not affected the initial exponential decay part of the glitch. The total fractional crustal superfluid moment of inertia involved in G7 is $I_{\rm cs}/I=2.05(1.98)\%$.
  • G8: We fit the eighth glitch (G8) of PSR J1420--6048 with Eqs (7), (10) and (14). As seen in Table 4, the exponential decay time-scale obtained from the vortex creep model application $\tau_{\rm creep,exp}=86(39)$ d is slightly different than $\tau_{\rm d}=100(6)$ d found from fit with Eq. (2) (see Table 3) again due to the initial exponential dependencies of Eqs. (10) and (14). The recoupling time-scale for the crustal superfluid associated with G8 is $\tau_{\rm s}=26(11)$ d and smallest among the glitches of PSR J1048--6048 analysed in this study. This value is consistent with its lower superfluid reservoir involvement in the glitch. The total fractional crustal superfluid moment of inertia involved in G8 is $I_{\rm cs}/I=1.23(2)\%$.
In addition, Eq. (15) estimates that the superfluid recoupling time-scale is $\tau_{\rm nl}=$(16–763) d for PSR J1420--6048. This prediction matches quite well with the range $[\tau_{\rm nl},\tau_{\rm s}]=[26(11)-779(554)]$ d obtained from the fit results given in Table 4.
4.2 PSR J1028--5819
Figure 2: Glitch in PSR J1028--5819. The descriptions of panels (a), (b), (c) and (d) are consistent with Fig. 1.
PSR J1028--5819 exhibits all the typical characteristics of a glitching pulsar, including a short period of $P\sim 91.4$ ms, a high period derivative of $\dot{P}\sim 16.1\times 10^{-15}$, a relatively small characteristic age of $\tau_{\rm c}$ $\sim 90$ kyr, and a large spin-down luminosity of $\dot{E}\sim 8.3\times 10^{35}$ erg s-1 (Keith et al., 2008). In 2009, Fermi-LAT detected gamma-ray pulsations from PSR J1028--5819 for the first time (Abdo et al., 2009). Later, Lower et al. (2021) reported its first glitch event at MJD $\sim 57881(14)$ using Parkes timing data, with a fractional glitch size of $\Delta\nu/\nu\sim 2296.5(5)\times 10^{-9}$.
[…]
According to existing studies, large glitches with exponential recovery are often accompanied by long-term linear recovery, which continues until the next glitch event occurs (Yu et al., 2013). The large glitch in PSR J1028--5819 follows this trend. As shown in panel (c) of Fig. 2, the blue line represents the fitted result for the linear decrease in $\dot{\nu}$. We found that the post-glitch slope ($\ddot{\nu}\sim 2.14(5)\times 10^{-23}$ s-3) is nearly three times larger than the pre-glitch slope ($\ddot{\nu}\sim 7.35(1)\times 10^{-24}$ s-3), indicating that the glitch caused a permanent change in the spin-down rate.
We fit this glitch with Eqs. (7) and (14), and the fit result is shown in Fig. 2. The total fractional crustal superfluid moment of inertia involved in the glitch is estimated to be $I_{\rm cs}/I=3.35(16)\%$, as seen in Table 5. The vortex creep model predicts that the time to the next glitch following G1 is $t_{\rm th}=4890(269)$ d. Since not all of the angular momentum reservoir within the crustal superfluid was depleted during G1, PSR J1028--5819 may experience another glitch due to a different pinned superfluid layer before the recovery from the previous event is fully complete. The fit result for the crustal superfluid’s recoupling timescale, $\tau_{\rm nl}=71$ d, is close to the lower end of the theoretical estimate range $\tau_{\rm nl}=$ 108–5113 d, as derived from Eq. (15).
Figure 3: Glitches in PSR J1709--4429. The descriptions of panels (a), (b), (c) and (d) are consistent with Fig. 1. In panel (d), the red dashed lines are the fitting result of the exponential recovery process of post-glitch 4, 5 and 6, and the number of fitting exponential are 2, 1 and 1, respectively.
Figure 4: PSR J1709--4429 timing residuals of glitch 4. The three panels represent the timing residuals for the spin-down model with 0, 1, and 2 exponential decay terms, respectively. The vertical gray dashed line and the number of parentheses at the top represent the glitch epoch and the reference number of glitch occurrences, respectively.
4.3 PSR J1709--4429 (B1706--44)
In 1992, the pulsed radio emission from PSR J1709--4429 was discovered for the first time in the Parkes high-frequency survey project (Johnston et al., 1992). Thompson et al. (1996) detected gamma-ray pulsations with EGRET, and Chandra found X-ray pulsed radiation in 2002 (Gotthelf et al., 2002). This is a bright Vela-like pulsar with a period of 102.5 ms, characteristic age of 17.5 kyr and the spin-down energy loss rate of $\dot{E}\sim 3.4\times 10^{36}$ erg s-1 (Wang et al., 2000). In 1995, the glitch event of PSR J1709--4429 was reported for the first time (Johnston et al., 1995), and five glitches were successively published in the following timing observations (Weltevrede et al., 2010; Yu et al., 2013; Lower et al., 2018, 2021).
[…]
For glitch 4, Yu et al. (2013) reported that post-glitch behaviour has only one exponential recovery. However, when we perform a fit combining Fermi-LAT and Parkes timing data, two exponential decay characteristics are revealed in the timing residuals, as shown in Fig. 4. The first exponential decay feature is $Q$1 = 0.0104(5) and $\tau_{\rm d1}$ = 72(4) d, and the second exponential decay feature is $Q$2 = 0.006(1) and $\tau_{\rm d2}$ = 4.2(6) d. Thus, the transient increment in the frequency and frequency first derivative is $\Delta\nu/\nu\sim 2777(4)\times 10^{-9}$ and $\Delta\dot{\nu}/\dot{\nu}\sim 64(15)\times 10^{-3}$, respectively.
For glitches 5 and 6, the glitch parameters obtained by our fitting are consistent with the previously reported results within the error range, and the detailed results are shown in Table 3. We performed exponential and linear fits to the evolution of $\dot{\nu}$, with the results displayed as red and blue dashed lines, respectively, in panel (c). For glitch 4, both single and double exponential functions were used to model the post-glitch exponential recovery of $\dot{\nu}$. The accuracy of fit was evaluated using the coefficient of determination ($R^{2}=1-\frac{\sum_{\rm i}(Y_{\rm i}-\overline{y})^{2}}{\sum_{\rm i}(y_{\rm i}-\overline{y})^{2}}$), where $y_{\rm i}$ represents the data points, $\overline{y}$ is their mean, and $Y_{\rm i}$ is the fitted value.
By fitting the exponential recovery of $\dot{\nu}$, we found that glitch 4 achieved the highest accuracy of fit ($R^{2}=0.95$) with a double exponential model, compared to $R^{2}=0.70$ for a single exponential fit. The result is consistent with that of the two exponential recovery procedures observed in the data processing. For glitches 5 and 6, an exponential term accurately describes the recovery of $\dot{\nu}$, with the fit achieving an accuracy of over 0.99. Interestingly, although the linear recovery slopes change significantly after glitches 1, 2, and 3 (Yu et al., 2013), our linear fitting results show that the linear recovery process after glitches 4, 5, and 6 exhibits a similar slope, which is approximately $\ddot{\nu}\sim 2.0\times 10^{-22}$ s-3.
We apply the vortex creep model to fit the three glitches, using Eqs. (7) and (14), and the fit result is shown in Fig. 3 panel (d):
  • G4: The exponential decay time-scale obtained from the vortex creep model application $\tau_{\rm creep,exp}=86(9)$ d is slightly different than $\tau_{\rm d}=72(4)$ d found from fit with Eq. (2) (see Table 3) again due to the initial exponential dependency of Eq. (14).  [^7] We mention that, alternatively, the $\tau_{\rm d}=4.2$ d time-scale exponential recovery associated with G4 of PSR J1709--4429 either arises from shallow densities of the crust (according to the vortex resistive model) or intermediate densities of the core region (according to the vortex cluster model). See more related discussions in Sec. 5.3. The vortex creep model prediction for the time to the next glitch from G4 is $t_{\rm th}=1680(16)$ d and in qualitative agreement with the observed inter-glitch time $t_{\rm obs}=1648(4)$ d. The total fractional crustal superfluid moment of inertia involved in G4 after corrected for the crustal entrainment effect is $I_{\rm cs}/I=2.30(2)\%$. The recoupling time-scale for the crustal superfluid associated with G4 is $\tau_{\rm nl}=3.4$ d.
  • G5: The exponential decay time-scale obtained from the vortex creep model application $\tau_{\rm creep,exp}=95(19)$ d is slightly different than $\tau_{\rm d}=61(6)$ d found from fit with Eq. (2) again due to the initial exponential dependency of Eq. (14). The vortex creep model prediction for the time to the next glitch from G5 is $t_{\rm th}=1861(25)$ d and matches with the observed inter-glitch time $t_{\rm obs}=1836(4)$ d quite well. The total fractional crustal superfluid moment of inertia involved in G5 is $I_{\rm cs}/I=2.20(3)\%$. The recoupling time-scale for the crustal superfluid associated with G4 is $\tau_{\rm nl}=28$ d.
  • G6: The exponential decay time-scale obtained from the vortex creep model application $\tau_{\rm creep,exp}=93(9)$ d differs from $\tau_{\rm d}=47(8)$ d obtained through fit with Eq. (2) again due to the initial exponential dependency of Eq. (14). The vortex creep model prediction for the time to the next glitch from G6 is $t_{\rm th}=1861(25)$ d, similar to the its previous glitch. The total fractional crustal superfluid moment of inertia involved in G6 is $I_{\rm cs}/I=1.85(2)\%$. The recoupling time-scale for the crustal superfluid associated with G6 is $\tau_{\rm nl}=3.2$ d.
4.4 PSR J1718--3825
Figure 5: Glitches in PSR J1718--3825. The descriptions of panels (a), (b), and (c) are consistent with Fig. 1. In Panel (c), the blue dashed curves represent the linear fitting of the linear recovery process at pre- and post-glitch. The purple curves show the fit with the vortex bending model (see Sec. 5).
PSR J1718--3825 is a young pulsar, first discovered by Manchester et al. (2001) in the radio band. It has a characteristic age of 89.5 kyr, a spin period of 74.7 ms, and a relatively high spin-down luminosity of $\dot{E}\sim 1.3\times 10^{36}$ erg s-1. In addition to an associated X-ray nebula (Hinton et al., 2007), it is also a source of gamma-ray emission (HESS J1718) within the supernova remnant RX J1713.7--3946 (Aharonian et al., 2007).
We analysed more than 14 yr of timing data for PSR J1718--3825 in the radio and gamma-ray bands. A total of 3 glitches were identified, of which glitch 1 and 2 have been reported in the radio band (Yu et al., 2013; Lower et al., 2021), and glitch 3 is a new small glitch. By examining changes in the timing residuals, it was deduced that glitch 3 occurred at MJD $\sim$ 59121(8). Using the fitting method, we determine the fractional glitch size to be $\Delta\nu/\nu\sim 1.9(2)\times 10^{-9}$ and the first frequency derivative to be $\Delta\dot{\nu}/\dot{\nu}\sim-0.12(9)\times 10^{-3}$. We also used the pre- and post-glitch timing data of glitch 2 to analyse the timing noise, as the time span exceeds three years. It was found that the spectral index before and after glitch 2 remained consistent at 6.
In Fig. 5, the three detected glitches in PSR J1718--3825 are presented. It is clear that $\dot{\nu}$ only experienced a long-term linear recovery after the glitches, and there may be stable $\ddot{\nu}$. Therefore, based on the glitch epochs, the $\dot{\nu}$ data were divided into four intervals and linear fits were applied to each interval to determine $\ddot{\nu}$. The fits revealed that, under the influence of glitch, $\ddot{\nu}$ has undergone obvious changes, with values of $2.0(3)\times 10^{-23}$ s-3, $2.887(3)\times 10^{-23}$ s-3, $1.65(3)\times 10^{-23}$ s-3, and $2.61(4)\times 10^{-23}$ s-3, respectively. These fitted values are represented by the blue dashed lines in panel (c) of Fig. 5. A comparison with predictions from the superfluid vortex bending model is also provided for the first (G1) and second (G2) glitch events. Detailed discussions are included below in Sec. 5.
4.5 PSR J1509--5850
PSR J1509--5850 is a middle-aged pulsar with a characteristic age of 154 kyr (Kramer et al., 2003). We have analysed nearly 14 yr of timing data in both the radio and gamma-ray bands and found no glitches within the observed data range. Our study of the spin-down rate evolution suggests that PSR J1509--5850 exhibits a roughly stable $\ddot{\nu}$, as shown in Fig. 6. Using the least squares method for linear fitting, we determined $\ddot{\nu}$ to be $1.328(4)\times 10^{-24}$ s-3, with the fitted result indicated by the blue dotted line in Fig. 6. Additionally, the figure presents a fit based on the superfluid vortex bending model, which describes well the $\dot{\nu}$’s oscillation (see Sec. 5).
Figure 6: Evolution of spin-down rate over time for PSR J1509--5850, along with a comparison between a linear fit (blue) and the vortex bending model (purple) describing well the oscillation, as discussed in Sec. 5.
5 DISCUSSION
5.1 Oscillations in the spin-down rate
Table 6: Fit results for the application of the vortex bending model. See text for details.
         Parameter          J1509--5850          J1718--3825 G1          J1718--3825 G2
         Data Range (MJD)          54303–59510          54997–57940          57964–59113
         $A$ (10-10 s-2)          6.86(41)          2.43(58)          10.95(26)
         $\tau$ (d)          5928          619090          218
         $\Omega_{0}$ (10-8 rad s-1)          2.63(27)          11.07(1.53)          19.94(56)
         $\phi_{\rm{b}}$ (rad)          0.20(53)          2.53(2.32)          1.50(2)
         $\dot{\nu}_{0}$ (10-12 s-2)          $-1.15926(1)$          $-2.36321(5)$          $-2.35852(2)$
         $\ddot{\nu}_{0}$ (10-24 s-3)          1.34(4)          20.25(32)          16.71(45)
         $T_{\rm osc}$ (yr)          7.567          1.799          0.998
         $\ell_{\rm v}$ (105 cm)          3.72          0.98          0.54
Commonly, the observed spin evolution of pulsars is simply attributed to the emission of magnetic dipole radiation and ejection of charge particles from the magnetosphere (pulsar wind; Manchester & Taylor 1977). Nevertheless, inside the pulsars, vortex lines could experience bending due to pinning in the inner crust and outer core, and applies torque on the surface of a neutron star as a result of their time variable coupling to the magnetospheres after transient events like mode switching, giant pulses, pulsar wind ejection, radiative bursts and glitches. Combining the linear creep response, the vortex bending results in an internal torque with vortex-bending-assisted oscillation modes and a damped sinusoidal-like oscillations in the star’s spin-down rate evolution (Gügercinoğlu et al., 2023).
After considering the negative braking index due to magnetic field change, the long-term spin-down properties are expressed as follows (Gügercinoğlu et al., 2023):
$\nu(t)=A\exp\left(-\frac{t}{2\tau}\right)\sin\left(\Omega_{0}t+\phi_{\rm{b}}\right)+\left(\dot{\nu}_{0}+\frac{1}{2}\ddot{\nu}_{0}t\right)t\ ,$ (18)
and
$\dot{\nu}(t)=$ $A\Omega_{0}\exp\left(-\frac{t}{2\tau}\right)\cos\left(\Omega_{0}t+\phi_{\rm{b}}\right)$ (19)
$-\frac{A}{2\tau}\exp\left(-\frac{t}{2\tau}\right)\sin\left(\Omega_{0}t+\phi_{\rm{b}}\right)$
$+\dot{\nu}_{0}+\ddot{\nu}_{0}t\ .$
Here, $A$, $\Omega_{0}$ and $\phi_{\rm{b}}$ are the amplitude, frequency and phase of the oscillations arising from the internal torque contribution due to vortex bending. $\dot{\nu}_{0}$ and $\ddot{\nu}_{0}$ are the spin-down rate and the second time derivative of the spin frequency of the neutron star due to external torques acting on its surface. The damping time of the oscillations in the spin-down rate $\tau$ is related to the linear creep recoupling time $\tau_{\rm l}$ given by Eq. (8) as $\tau=(I_{\rm c}/I)\tau_{\rm l}$. Here, $I$ is the total moment of inertia of the neutron star. $I_{\rm c}$ is the sum of the moment of inertia of the neutron star normal matter and core superfluid/superconducting components that couples to its surface on short timescales and tied by magnetic field crossing the whole star. Basically, $I_{\rm c}$ excludes the inner crust superfluid and the toroidal field region in the outer core of neutron stars. Oscillation time is defined as
$T_{\rm osc}=\frac{2\pi}{\Omega_{0}}=\frac{2\pi}{\omega_{0}}\left[1-\left(\frac{1}{2\tau\omega_{0}}\right)^{2}\right]^{-1/2}\ ,$ (20)
with
$\omega_{0}=\left(\frac{I_{\rm b}I}{I_{\rm cs}^{\prime}I_{\rm c}}\frac{10\Omega_{\rm c}\kappa}{\pi\ell^{2}}\right)^{1/2}\ .$ (21)
Here, $I_{\rm cs}^{\prime}$ is the moment of inertia of the crustal superfluid, $I_{\rm b}$ is the moment of inertia of the regions in which vortex lines has bent over due to pinning (crustal superfluid plus toroidal field region in the outer core) with $\ell$ being the length-scale over which vortices bent and $\kappa$ is the vorticity quantum attached to each line.
We apply the vortex bending model to the timing solution of PSR J1509--5850 between MJD 54303–59510, see Fig. 6. The model free parameters corresponding to the vortex bending model’s application are given in Table 6. If we attribute the large frequency second time derivative of $\ddot{\nu}\cong 1.37\times 10^{-24}$  s-3 to an unresolved Vela-like glitch occurred before the time interval we considered, then Eq. (24) with the typical value $\Delta\dot{\nu}_{\rm p}/\dot{\nu}=5\times 10^{-3}$ gives the estimate of $t_{\rm ig}=158$ yr for the time between two successive giant glitches hypothesised to be experienced by this pulsar. Thus, we predict that PSR J1509--5850 will glitch again within the same timescale. Future observations of this pulsar are required to test our assertion.
We also analyse the timing data of PSR J1718--3825 between MJD 54303–59491. Since there is not enough data for the third glitch (G3), we made an application of the vortex bending model to its first (G1) and second (G2) glitch events. See Fig. 5, panel (c). The fit parameters corresponding to the vortex bending model are given in Table 6. While the vortex bending model slightly improves the fit to the data of G1 compared to the linear one, it provides an excellent fit to the G2 data reducing the residuals significantly over the linear one. In particular, the damping timescale of the oscillations in the spin-down rate is well constrained to be $\tau=218$ d. When assessed with Eq. (8), this would provide important aspects for the microphysical parameters related to nuclear pinning mechanism as well as temperature estimate for the neutron star crust. Also, the length-scale $\ell$ found from the fits would provide a new method to put constraints on the crustal thickness and in turn on the EOS of the neutron stars. It should be also important to address these constraints in light of the sinusoidal-like oscillations observed in the spin-down rates of various pulsars.
5.2 Inter-glitch Braking Index
Table 7: The braking indices $n_{\rm{ig}}$ predicted in the vortex model. Also shown is the comparison of the observed inter-glitch $\ddot{\nu}_{\rm obs}$ with the vortex creep model prediction $\ddot{\nu}_{\rm ig}$ given by Eq. (23).
PSR Gl. No. Epoch $n_{\rm{ig}}$ $\ddot{\nu}_{\rm ig}$ $\ddot{\nu}_{\rm obs}$
(MJD) (10-24s-3) (10-24s-3)
J1028--5819 1 57904(6) 67(4) 22.9(13) 20(3)
J1420--6048 4 54653(19) 96.8(26) 2097(56) 1145(26)
5 55400(9) 59(6) 1281(139) 932(36)
6 56256(1) 44.2(23) 958(50) 1172(59)
7 57216(12) 30.6(9) 664(20) 779(9)
8 58555(2) 54.1(11) 1170(24) 551(11)
J1709--4429 4 54691(2) 42.8(5) 343(4) 351(4)
5 56339(2) 23(4) 188(31) 135(91)
6 58175(2) 35(4) 282(30) 227(43)
If only the external torque is considered, based on the use of torque laws of the form $-K\Omega^{n}$, where $K$ is a constant and the braking index $n={\nu\ddot{\nu}}/{\dot{\nu}}^{2}$ is assumed to be 3 as predicted by the magnetic dipole model (Manchester & Taylor, 1977). Following the vortex creep model (Alpar et al., 1984; Gügercinoǧlu & Alpar, 2017), glitches may occur as a result of collective unpinning of a very large number of vortex lines in the crustal superfluid and inter-glitch spin evolution of a pulsar is dominated by the response of the star to the variations in the superfluid internal torques’ coupling due to such vortex discharges. In terms of the model parameters theoretical expectation of the inter-glitch braking index can be recast as (Alpar & Baykal, 2006):
$n_{\rm{ig}}=(\beta+1/2)\left[\left(\Delta\dot{\nu}_{\rm p}/\dot{\nu}\right)_{-3}^{2}/(\Delta\nu_{\rm p}/\nu)_{-6}\right]\ ,$ (22)
where the glitch magnitudes in the rotation and the spin-down rates are expressed in units of $10^{-6}$ and $10^{-3}$, respectively. Also, we have used the shorthand notation $\beta=I_{\rm B}/I_{\rm A}$. As a fiducial value, in our calculations below we take $\beta=4.11$, which was obtained for PSR J1048–5832 (Liu et al., 2024). In comparison, the $constant$ braking index ($n$) before and after the glitch were estimated according to the fitted $\ddot{\nu}$ and the parameters ${\nu}$, $\dot{\nu}$ in Table 2. We observed stable values of $\ddot{\nu}$ for PSRs J1028--5819, J1509--5850 and J1718--3825 and fitted the slopes of $\ddot{\nu}$ for each. In the case of PSR J1028--5819, the braking indices before and after the glitch are 21.66(4) and 62.8(4), respectively. For glitch 2 of PSR J1718--3825, the braking indices before and after the glitch are 69.35(7) and 41.0(8), respectively. Although we did not identify a glitch event for PSR J1509--5850, its $\dot{\nu}$ has steadily evolved over 14 yr, resulting in a braking index of $n=11.12(2)$. These non-canonical braking indices may be indicative of internal torque at work, likely associated with the dynamics of nucleon superfluidity in this context.
Moreover, the most prominent effect of vortex unpinning on the spin parameters is $\ddot{\nu}$ and given by
$\ddot{\nu}_{\rm ig}=\left(\beta+\frac{1}{2}\right)\left(\frac{\Delta\dot{\nu}_{\rm p}}{\dot{\nu}}\right)^{2}_{-3}\left(\frac{\Delta\nu_{\rm p}}{\nu}\right)^{-1}_{-6}\left(\frac{\dot{\nu}^{2}}{\nu}\right)\ .$ (23)
The inter-glitch time is expressed as
$t_{\rm ig}=\frac{\Delta\dot{\nu}_{\rm p}}{\ddot{\nu}_{\rm ig}}=2\times 10^{-3}\left(\beta+\frac{1}{2}\right)\tau_{\rm sd}\left(\frac{\Delta\dot{\nu}_{\rm p}}{\dot{\nu}}\right)^{-1}_{-3}\left(\frac{\Delta\nu_{\rm p}}{\nu}\right)_{-6}\ .$ (24)
Comparison of the prediction of the vortex creep model [Eq. (23)] with observed inter-glitch $\ddot{\nu}$ for our sample of glitches is given in Table 7. An inspection to the Table 7 reveals that the agreement between the vortex creep model predictions and observations are qualitatively good except for G8 of PDR J1420--6048. The reason for the discrepancy may be a persistent step increase that not resolved which leads to a somewhat larger estimate in $\Delta\dot{\nu}_{\rm p}$. Future observations will help to test this hypothesis.
5.3 Exponential Recovery Mechanisms
Figure 7: Exponential decay time-scale of the the neutron star core predicted by the vortex cluster model [Eq. (25)] (in logarithmic scale) versus matter density for PSR J1028--5819 (blue dotted curve), PSR J1420--6048 (green solid curve) and PSR J1709--4429 (red dashed curve).
For glitch events with large frequency changes, exponential recovery may occur after glitch (Yu et al., 2013; Lower et al., 2021; Basu et al., 2022). The exponential recovery timescale is generally between 10 and 300 d, with a longer timescale for older pulsars (Yu et al., 2013). There may be more than one exponential recovery process after the glitches, for example, Vela pulsar has been detected four exponential recovery processes in its glitch 14 (Yu et al., 2013; Dodson et al., 2002). In the vortex creep model discussed above, the outer core superfluid’s response to a glitch involves exponential relaxation through vortex creep across magnetic flux tubes. The corresponding decay timescale has been calculated for several gamma-ray glitching pulsars (Gügercinoğlu et al., 2022; Liu et al., 2024). Considering the complex dynamical behaviour observed in the data it would be worthwhile to explore alternatives that differ from the vortex creep model in several important details. Here, we consider two alternative mechanisms that could explain the post-glitch exponential recovery and apply them to the observational data for the sample analysed in this study.
Sedrakian & Sedrakian (1995) considered a different mechanism based on vortex line-flux tube interaction originated from the whole neutron star core. According to them, initial magnetic field of a neutron star inherited from its progenitor star is expelled out from the core region as a result of setting up of Meissner currents after transition to superconducting state. However, up to a certain distance from its centre known as the London penetration depth, each vortex line is attached with magnetic flux due to proton supercurrent circulating around them. Immediate vicinity of a vortex line is surrounded by more numerous flux tubes and a so-called vortex cluster is formed. Electrons’ scattering off such vortex clusters couples the superfluid/superconducting core of a neutron star to its crust and leads to exponential decay following glitches. Vortex cluster model predicts for the decay time-scale (Sedrakian & Sedrakian, 1995),
$\tau_{\rm d,cluster}=\frac{\sqrt{3}\pi}{2048}\frac{\hbar}{m_{\rm e}}\frac{\rho_{\rm e}}{\rho_{\rm s}^{*}}\frac{1}{k_{\rm e}\lambda_{\rm p}^{3}}\mu|k|\left(\frac{\xi_{\rm p}}{\lambda_{\rm p}}\right)^{\frac{1}{|k|}}P^{2}\ ,$ (25)
where $\hbar$ is the Planck constant divided by $2\pi$, $m_{\rm e}$ is the mass of the electron, $\rho_{\rm e}$ is the mass density of electrons, $\rho_{\rm s}^{*}$ is the mass density of the entrained superfluid neutrons, $k_{\rm e}$ is the electron wave number, $\lambda_{\rm p}$ is the London penetration depth for proton supercurrents, $\xi_{\rm p}$ is the coherence length for superconducting protons, $k\equiv(m_{\rm p}^{*}-m_{\rm p})/m_{\rm p}$ is the entrainment coefficient for the superconducting protons, $\mu$ is the magnetic permeability of vortex cluster and $P$ is the spin period of the pulsar. For calculating these parameters Sedrakian & Sedrakian (1995) employed moderately stiff AV14+UV EOS of Wiringa et al. (1988), and took microphysical quantities associated with superfluidity/superconductivity from Baldo et al. (1992). We adopted these values as given in tables 1 and 2 of their paper. In Fig. 7 we give the exponential decay timescale estimate Eq. (25) for the three glitching gamma-ray pulsars studied in this paper. Due to large variation of the model free parameters throughout of the core region, estimate Eq. (25) ranges from several minutes to decades for a given pulsar and in practice able to explain all exponential recovery time-scales seen across all glitching pulsar population. And, as the only parameter differing among different pulsars in their model is the pulsar spin period, which occupies a narrow parameter space, the curves in Fig. 7 are quite similar.
Figure 8: Exponential decay time-scale of the neutron star crust predicted by the vortex resistive model [Eq. (26)] (in logarithmic scale) versus matter density for PSR J1028--5819 (blue dotted curve), PSR J1420--6048 (green solid curve) and PSR J1709--4429 (red dashed curve).
Jones (1990) considered the energy transfer process associated with the superfluid vortex line past the crustal lattice and mutual friction coupling to the phonons. At relatively low velocity lag between the vortex and the lattice the interaction is mediated by excitation and propagation of sound waves. Jones (1993) argued that since the neutron star crustal lattice is described by a polycrystalline structure and vortex lines have different orientations with respect to the nuclear lattice, the pinning forces arising from the randomly distributed point defects of lattice around the axis of vortex may largely cancel each other, at least for a portion of the inner crust superfluid. So, the dominant force acting on a vortex moving through the inner crust is the resistive force due to energy carried over to the phonons by above mentioned mechanism. Such coupling leads to exponential decay when vortices unpinned and displaced with respect to the ionic crystal. Decay timescale is given by (Jones, 1990)
$\tau_{\rm Jones90}=\frac{1}{2\mathcal{B}\Omega}=\frac{16\pi^{3/2}}{3}\frac{\xi^{3}_{\rm n}}{a}\frac{Mc^{3}_{\rm s}}{E^{2}_{\rm p}}\frac{\kappa\rho_{\rm s}}{\Omega},$ (26)
where, $a$ is the lattice constant, $\xi_{\rm n}$ is the vortex core radius or neutron superfluid’s coherence length, $M$ is the total mass in atomic nuclei, $E_{\rm p}$ is the pinning energy between a nucleus and vortex, $\kappa$ is the vorticity, $\rho_{\rm s}$ is the superfluid neutron’s mass density and $\Omega=2\pi/P$ is the angular velocity of the neutron star. In the above expression, $c_{\rm s}$ is the lattice phonon velocity and defined as the ratio of the plasma frequency to the Debye wave number. On the other hand, $\mathcal{B}$ is the mutual friction parameter and allows for comparison of the strength of this mechanism with other processes prevailing inside neutron stars (Haskell et al., 2012). Microphysical parameters in Eq. (26) were calculated for five distinct densities in the crust considered in Negele & Vautherin (1973) and were given in table 1 of Jones (1990). In Fig. 8 we give the exponential decay timescale estimate Eq. (26) for the three glitching gamma-ray pulsars studied in this paper. Although Eq. (26) does not exhibit diversity as much as that predicted by Eq. (25) for relevant neutron star parameters, the exponential decay time for this process also varies greatly by three orders of magnitude. Again, the curves in Fig. 8 are very similar to each other because the spin period, which is the only variable quantity other than nuclear physics parameters, does not change much among pulsars considered in this study. Note that, the sharp changes of lines in the figure do not correspond to physical discontinuities, but are due to a low number of data points (only 5 for each curve).
6 CONCLUSIONS
In this study, we combined the Fermi-LAT and the Parkes timing data of gamma-ray pulsars J1028--5819, J1420--6048, J1509--5850, J1709--4429 and J1718--3825 for more than 14 yr to study glitches. Our analysis has led to the identification of a total of 12 glitch events in four pulsars, among which there is one new glitch event. These glitches exhibit fractional sizes $\Delta\nu/\nu$ ranging from $1.9(2)\times 10^{-9}$ to $2964(4)\times 10^{-9}$. Furthermore, we have undertaken a comprehensive update of glitch-related parameters for glitches previously reported. An intriguing aspect of our study is the effective combination of the ToAs from both Fermi-LAT and Parkes, which not only augments the effective observation cadence of these pulsars, but more importantly, have the potential to fill the observational gap after the glitch occurs, thereby accurately revealing the exponential recovery process on the short time-scale of post-glitch. As a result, our investigation has unveiled several novel phenomena within the glitch recovery process.
In the case of glitch events 4, 6, and 8 of PSR J1420--6048, we discovered two linear recovery processes in the post-glitch evolution of $\dot{\nu}$. Furthermore, regarding glitch 8, we noted the presence of an exponential recovery process with $Q=0.0131(5)$, $\tau_{\rm d}=100(6)$ d. For PSR J1709--4429, we detected two exponential decay processes within glitch 4, with $Q1=0.0104(5)$, $\tau_{\rm d1}=72(4)$ d, $Q2=0.006(1)$, $\tau_{\rm d2}=4.2(6)$ d, and the parameter $\Delta\dot{\nu}/\dot{\nu}$ exhibited significant difference, standing at $64(15)\times 10^{-3}$. Meanwhile, we found that $\ddot{\nu}$ changed significantly during post-glitch linear recovery in glitches 1, 2 and 3, but not in glitches 4, 5 and 6. For PSR J1718--3825, a new small glitch was identified in MJD $\sim$ 59121(8), with glitch parameters are $\Delta\nu/\nu\sim 1.9(2)\times 10^{-9}$ and $\Delta\dot{\nu}/\dot{\nu}\sim-0.12(9)\times 10^{-3}$.
We connected the observed glitches with the dynamics of neutron vortex arrays inside the pulsars and discussed the theoretical interpretation of these events in the framework of vortex creep model. For our sample of glitches, application of the vortex creep model revealed that the total fractional moment of inertia of the crustal superfluid required to tapped at glitches is $\lesssim 3.5\%$, indicating that only crustal superfluid and probably some part of the outer core region of a neutron star are involved in glitches. Yet another implication of the application of the vortex creep model is that the model estimate for the time to the next glitch in general agrees well with the observed inter-glitch time for a given pulsar’s particular glitch. Thus, by doing timing fit to a particular glitch future observations can be planned for corresponding pulsar to establish any variations in the pulse profile concurrent with the glitch occurrence. Such a detection will enable us to connect the magnetospheric physics to the interior dynamics and disclose the interplay between them. Different parts of the post-glitch recovery, in particular the exponential relaxation can be studied to infer microphysics related to the neutron star crustal and core matter. We plan to compute the relaxation time derived from microscopic two- and three-body nuclear forces, incorporating pasta phases and proton effective masses self-consistently, as well as application to the observed cases in a separate paper. We’d also like to address the effects of more updated superfluidity/superconductivity-related interaction parameters by taking proximity etc. into account on the exponential recovery time estimate. Simulating pulsar glitches through experimental observations in rotating ultracold atoms offers exciting potential to constrain the complex physics associated with superfluid vortices (Poli et al., 2023). Future astrophysical observations will provide a larger sample, particularly with variations in post-glitch behaviour, promising greater precision in glitch constraints and offering insights into the dynamics of superfluid vortices beyond the coarse-grained two-component scenario. Additionally, pulsar glitches may imprint themselves on gravitational wave radiation, establishing a promising link between multi-messenger studies of these phenomena and insights into the dense matter EOS (Cheunchitra et al., 2024; Wilson & Ho, 2024; Yim et al., 2024).
Acknowledgements
The work is supported by National SKA Program of China (no. 2020SKA0120300), the Strategic Priority Research Program of the Chinese Academy of Sciences (grant no. XDB0550300), the National Natural Science Foundation of China (grant nos 12273028, 12041304 and 12288102), and the Major Science and Technology Program of Xinjiang Uygur Autonomous Region (grant no. 2022A03013, 2022A03013-4). EG acknowledges support from the National Natural Science Foundation of China (NSFC) programme 11988101 under the foreign talents grant QN2023061004L. ZZ is supported by the Natural Science Basic Research Program of Shaanxi (program no. 2024JC-YBQN-0036). SD is supported by the Guizhou Provincial Science and Technology Foundation (no. ZK[2022]304) and the Scientific Research Project of the Guizhou Provincial Education (no. KY[2022]132). PW acknowledges support from the National Natural Science Foundation of China under grant U2031117 and the Youth Innovation Promotion Association CAS (id. 2021055). DL is supported by the 2020 project of Xinjiang Uygur autonomous region of China for flexibly fetching in upscale talents. We acknowledge the use of the public data from the Fermi-LAT data archive. The Parkes radio telescope is part of the Australia Telescope National Facility which is funded by the Commonwealth of Australia for operation as a National Facility managed by CSIRO. This paper includes archived data obtained through the CSIRO Data Access Portal.
Data Availability
The data underlying this article will be shared on reasonable request to the corresponding authors.